OpenStudy (sloppycanada):

If y varies directly as x, and y = 15 as x = -3, find y for the x-value of 4. The answer is 2 If y varies inversely as x, and y = 5 as x = 6, find y for the x-value of 10. k = 60?

OpenStudy (sloppycanada):

@Loser66 Just making sure these answers are correct.

OpenStudy (loser66):

Nope. y =15, x =-3, hence k =-5, your equation is y = -5x, then if x =4, y = -20, not 2

OpenStudy (loser66):

if y is varies inversely as x, that is y = k/x , k/x = 5 when x =6, hence k = 30 , not 60 y = 30/x, if x =10 , y = 3 hey!! I worked!!! not you. why???

OpenStudy (sloppycanada):

I honestly didn't think I could plug in the new numbers into the first solution..

OpenStudy (loser66):

\(y =kx\\15 = k*(-3)\\k = \dfrac{15}{-3}=-5\)

OpenStudy (loser66):

hence \(y = (-5)x\)

OpenStudy (loser66):

Now, find y if x =4, just replace x =4 into it \(y = (-5)*4= -20\)